What is the effect on the volume of 1 mole of an ideal gas if both the pressure and the temperature are reduced to one- third of their initial values? Explain.
Answer #1

ANSWER

The effect of temperature and pressure on the volume of ideal gas can be estimated by ideal gas equation:

PV = nRT

P1V1 / T1 = P2V2 / T2

P1, V1 and T1 = initial pressure , volume and temperature respectively

P2 , V2 and T2  = final pressure , volume and temperature respectively

V2 = (P1V1 T2 / P2 T1 ) =======1)

P2 = P1 / 3 and T2 = T1 / 3

Substituting these values in equation (1)

V2 = (3 P1V1 T1 / 3 P1  T1 ) = V1

Hence final volume = initial volume. Thus there is no change in the volume of the gas.